The length of a rectangle is 3 in. more than twice its width. If the perimeter of the rectangle is 36 in., find the width of the rectangle.
Answer by ptaylor(2198)

Let x=width
Then 2x+3=length
Perimeter(P)=2L+2W ---So our equation to solve is:
36=2(2x+3)+2x get rid of parens
36=4x+6+2x collect like terms
36=6x+6 subtract 6 from both sides
6x+6-6=36-6
6x=30 divide both sides by 6
x=5 in------------------------------------width
2x+3=2*5+3=13 in-------------------------------length
CK
36=2*5+2*13
36=10+26
36=36
Hope this helps----ptaylor